# Polyhedral Collision Detection via Vertex Enumeration

### v1

@@ I INTRODUCTION
Handling collision detection in trajectory planning is hard, primarily because the degree to which two objects intersect in physical space (the signed distance) generally can not be represented by a continuous and differentiable function, as is typically required by solvers of constrained optimization problems. To the best of our knowledge, only the distance between spheres can be expressed in this way. For example, two circles in the plane with centers $c_{1}$ and $c_{2}$, and radii $r_{1}$ and $r_{2}$ are intersecting if and only if $\|c_{1}-c_{2}\|\leq r_{1}+r_{2}$. The situation is not as simple if the circles are replaced with a triangle and a rectangle. The most common way to circumvent this issue is by approximating exact collision detection, representing the relevant geometries as union of spheres [1, 2, 3, 4, 5, 6, 7, 8, 9]. This approach can lead to excessive conservatism for poor approximations, or accurately modeling the geometry can be computationally expensive.
@@
(ii) Attempting to wrap the low-level optimization collision detection problem as a continuous and differentiable function also faces significant difficulty, because most approaches invoke the implicit function theorem [13, 14, 15] to produce derivatives of the low-level optimal decision variables with respect to the high-level decision variables, while high-level decision variables appear as problem data (parameters) in the low-level. This approach is fundamentally flawed, because when inequality constraints are weakly active at the solution of the low-level problem, the implicit function theorem cannot be invoked. Derivatives of the optimal decision variables and objective value with respect to the problem data at these points generally do not exist. The set of points in problem data space where this occurs is normally of measure zero, so it is tempting to ignore any issues of non-differentiability. However, the points at which a low-level problem solution will fail to be differentiable are frequently encountered in bilevel collision detection schemes.
@@
  • A procedure and code repository for replacing nonsmooth constraints with multiple smooth constraints by enumerating the values obtained at all vertices of the feasible region of signed distance between polyhedra.
@@ II-A Problem Formulation
representing second-order rigid-body dynamics with full degrees of freedom (DoF), where $x_{t}\in\mathbb{R}^{6(d-1)}$ is the state at time $t$ in $d\in\{2,3\}$-dimensional space with full-DoF second-order controls $u_{t}\in\mathbb{R}^{3(d-1)}$. The initial state and control of the ego is $x_{0}$ and $u_{0}$. We use full state and control parametrization for nonlinear optimization, the decision variable is defined as $z_{t}=\begin{bmatrix}x_{t}&u_{t}\end{bmatrix}^{\intercal}$ for all $t\in\{1,\dots,T\}$. The complete decision variable $z=\{z_{1},\dots,z_{T}\}\in\mathbb{R}^{n}$ represents the whole trajectory, where $n=T\times 9(d-1)$. This decision variable is subject to Eq. 1, and other constraints such as saturation of controls, and collision detection constraints.
@@ II-B Convex Object Processing
Intuitively, the $\alpha$ is a scaling factor that uniformly inflates (or deflates) the convex object. The object inflates with around the point $c$ when $\alpha>0$, covering the entire space as $\alpha\to\infty$. When $\alpha=0$, Eq. 5 becomes the same as Eq. 4. When $\alpha<0$, the object deflates, becoming a singleton $\{c\}$ at $\alpha=-1$.
@@ II-C Signed Distance
$\mathrm{sd}(x_{t}):=\min_{\alpha\in\mathbb{R}}$ $\alpha$ (7)
$\mathrm{s.t.}$ $p\in\mathcal{O}_{1}^{x_{t}}(\alpha)\cap\mathcal{O}_{2}(\alpha).$
@@
Non-differentiability: For a $\{3,4\}$-dimensional polyhedron defined by $m_{\mathrm{sd}}$ inequality constraints, a vertex $w$ requires at least $\{3,4\}$ inequalities to become equalities. Every vertex can be attained by solving a linear system $Aw+b=0$, where $A\in\mathbb{R}^{(d+1)\times(d+1)}$, $b\in\mathbb{R}^{d+1}$. $A$ and $b$ depend on which inequalities we select to become equalities. If we traverse all possible selections, we can get all possible vertices.
@@
Given $x_{t}$, if there are only $d+1$ active constraints, there is only one optimal assignment, and we can get a unique assignment point $w$, so $\alpha$ is also unique, and the signed distance function $sd(x_{t})$ is smooth. However, when there are more than ${d+1}$ active constraints, the optimal assignment is not unique, and $sd(x_{t})$ is nonsmooth, because although $\alpha$ corresponding to these optimal assignments have the same value, they have different expressions and therefore different derivatives.
In Fig. 1, we show an example with two polygons and the corresponding 3-dimensional polyhedral LP space. There exist five feasible assignments, but only two values of $\alpha\in\{\alpha*,\alpha_{1}\}$. Clearly, $\alpha^{*}$ is the minimizer of Eq. 9, i.e. the actual signed distance. When $\alpha=\alpha^{*}$, the interception point $p$ can be any point on the segment corresponding to an edge in the 3-dimensional space. Any point on this edge is a solution to the LP. If the ego rotates slightly, this edge will degenerate to either one of the vertices that correspond to $\alpha=\alpha^{*}$. The other near feasible assignment of constraints correspond to $\alpha=\alpha_{1}$, which is also marked in Fig. 1.
@@ III Collision Detection via Vertex Enumeration
To overcome this, we propose slots, placeholders which are dynamically filled with appropriate vertex values. Specifically, we let $N$ be the maximum number of vertices we wish to consider at any time. At each value of $x_{t}$, we enumerate all assignments, discard infeasible ones, and then sort the remaining $K$ assignments by $\mathrm{sd}_{k}(x_{t})$ value, and place $\mathrm{sd}_{k}(x_{t})$ expressions of the $N$ smallest ones into the slots as our signed distance constraints. If $K<N$, we copy the $\mathrm{sd}_{k}(x_{t})$ expression with the largest $sd_{k}(x_{t})$ value to fill the remaining slots.
The slot concept we introduce does not completely address the issue at hand. As $x_{t}$ is updated in the optimization, the particular vertices which fill each slot can change. More specifically, since each vertex is an assignment point, the assignment points which fill each slot can change. Although the assignment points are themselves smooth functions of $x_{t}$, this implies that the output value of each slot is not a smooth function of $x_{t}$. Nevertheless, we propose simply ignoring this fact when computing the derivative of each constraint slot (as needed by optimization solvers), and simply treating the derivative of the assignment point value (for the given value of $x_{t}$) as the derivative of the constraint slot.
The reformulation we have so far presented may seem counterintuitive. Initially, the signed-distance constraint appearing in Eq. 2 was a continuous but non-differentiable constraint. Now, using the slot formulation of Eq. 10, there are $K$ constraints which are also continuous but non-differentiable. The benefits to the reformulation emerge from two key aspects.
The first benefit of this new formulation is representing derivative information at points of non-differentiability. In the formulation Eq. 10, non-differentiability occurs whenever there are multiple assignment points which could be placed into the constraint slots due to equivalent value. The only constraint slots which can ever be “active” are the first slot, and any other slots who are tied in value at a particular iteration. After some perturbation to $x_{t}$, the values will diverge and only the smaller of the two will remain relevant for the purposes of the optimization Eq. 10. It is plausible that the derivative information for that slot was incorrect before the perturbation, but if that was the case, then the correct derivative information must have been associated to one of the other tied slots. This enables the solver to essentially capture the subdifferential of the signed distance at these points of non-differentiability.
@@ IV-A Baselines
The first baseline method we compare to is the separating hyperplanes method [19, 20]. It is a common and efficient approach that utilizes the hyperplane separation theorem, which states there exists a separating hyperplane that separates two nonempty nonintersecting closed polyhedral objects. It is very fast for simple settings, where there is one ego and one obstacle.
The second baseline method is [15], where the authors formulate the collision detection problem using a signed distance function as a convex optimization problem with conic constraints. However, this approach is flawed, because when inequality constraints are weakly active at the solution of the low-level problem, the implicit function theorem cannot be evoked, which our formulation avoids.
We compare our method to the baselines in terms of speed and reliability. Our metric for speed is the mean MCP solve time as solved by the PATH Solver. The reliability metric is more tricky, we provide two metrics to represent reliability: (i) Success rate in terms of successful MCP solution achieved by PATH Solver represents the reliability of the method in terms success converging to a local optimum, (ii) and the mean cost of successes indicates how reliably the formulation achieves a low-cost local optimum. We define the success rate as the ratio of successful solutions to the total number of samples.
@@ IV-C Results
In this section we discuss large-scale 2-dimensional simulation results for our method and baselines on randomized obstacles and initial conditions for different problems. As we will show in this section, our method achieves higher success rates in settings where the ego needs to navigate through and pack into multiple obstacles, but it is slower than separating hyperplane formulation. In our experiments, we compare our method against two baseline methods for six different problems, and we randomly generate initial conditions and the obstacles (if applicable).
@@
We showcase the success rate of our method compared to baselines in Table I, where we use boldface to indicate the highest success rates. We see that the separating hyperplane approach is less reliable locally than our method for all but the simplest problem, simple packing. Our method achieves greater than $90\%$ success for all problems, beating the baselines in all settings except for the piano problem. We show the resulting cost for the successful solutions in Table II (failed MCP solutions are not included in this metric), where we use boldface to indicate the best cost. Table II tells us how successful each formulation is in terms of global success, such as being able to pass through a gap to reach the goal position. Our method outperforms the baselines in packing type problems, which means it packs more reliably compared to the baselines. In gap type problems however, our method does not pass through the hole as often as [15] does.
@@
In this section we demonstrated that our proposed method, nonsmooth signed distance enumeration, achieves superior accuracy solving complex collision detection problems, while remaining competitive in terms of performance. We have demonstrated that our formulation is more reliable compared to the separating hyperplane formulation, however it is slower. This is not surprising, as separating hyperplane formulation is quite efficient provided with a good initialization. Furthermore, we demonstrated that our method is more reliable than a comparable but flawed approach in the literature for the problem of polygon collision detection, especially for packing problems where incorrect derivatives have a higher chance of becoming a problem for the solver.
@@ V CONCLUSION
In this paper we proposed a procedure for dealing with nonsmooth constraints by enumerating the values obtained at all vertices of the feasible region of signed distance between polyhedra. We then demonstrated our formulation on several challenging collision detection problems in 2 and 3-dimensional settings, and compared them to common existing formulations. We found out that our method outperforms common baselines in terms of success rate. We found that our formulation is slower than separating hyperplanes but it is more reliable.

### v2

@@ I INTRODUCTION
Handling collision detection in trajectory planning is hard, primarily because the degree to which two objects intersect in physical space (the signed distance) generally can not be represented by a continuous and differentiable function, as is typically required by solvers of constrained optimization problems. To the best of our knowledge, only the distance between spheres can be expressed in this way. For example, two circles in the plane with centers $c_{1}$ and $c_{2}$, and radii $r_{1}$ and $r_{2}$ are intersecting if and only if $\|c_{1}-c_{2}\|\leq r_{1}+r_{2}$. The situation is not as simple if the circles are replaced with a triangle and a rectangle. The most common way to circumvent this issue is by approximating exact collision detection, representing the relevant geometries as union of spheres [1, 2, 3, 4, 5, 6, 7, 8, 9]. This approach can lead to excessive conservatism for poor approximations, or accurately modeling the geometry can be expensive.
@@
(ii) Attempting to wrap the low-level optimization collision detection problem as a continuous and differentiable function also faces significant difficulty, because most approaches invoke the implicit function theorem [13, 14, 15, 16] to produce derivatives of the low-level optimal decision variables with respect to the high-level decision variables, while high-level decision variables appear as problem data (parameters) in the low-level. This approach is fundamentally flawed, because when inequality constraints are weakly active at the solution of the low-level problem, the implicit function theorem cannot be invoked. Derivatives of the optimal decision variables and objective value with respect to the problem data at these points generally do not exist. The set of points in problem data space where this occurs is normally of measure zero, so it is tempting to ignore any issues of non-differentiability. However, the points at which a low-level solution will not be differentiable are frequently encountered in bilevel collision detection.
@@
  • A procedure and code repository [^1] The code is available at https://github.com/VAMPIR-Lab/PolyPlanning. for replacing nonsmooth constraints with multiple smooth constraints by enumerating the values obtained at all vertices of the feasible region of signed distance between polyhedra.
@@ II-A Problem Formulation
representing second-order rigid-body dynamics with full degrees of freedom (DoF), where $x_{t}\in\mathbb{R}^{6(d-1)}$ is the state at time $t$ in $d\in\{2,3\}$-dimensional space with full-DoF second-order controls $u_{t}\in\mathbb{R}^{3(d-1)}$. The initial state and control of the ego is $x_{0}$ and $u_{0}$. We use full state and control parametrization for nonlinear optimization, the decision variable is defined as $z_{t}=\begin{bmatrix}x_{t}&u_{t}\end{bmatrix}^{\intercal}$ for all $t\in\{1,\dots,T\}$. The complete decision variable $z=\{z_{1},\dots,z_{T}\}\in\mathbb{R}^{n}$ represents the whole trajectory, where $n=T\times 9(d-1)$. This decision variable is subject to Eq. 1, and other constraints such as saturation of controls, and collision constraints.
@@ II-B Convex Object Processing
Intuitively, the $\alpha$ is a scaling factor that uniformly inflates (or deflates) the convex object. The object inflates with around the point $c$ when $\alpha>0$, covering the entire space as $\alpha\to\infty$. When $\alpha=0$, Eq. 5 is the same as Eq. 4. When $\alpha<0$, the object deflates to a singleton $\{c\}$ at $\alpha=-1$.
@@ II-C Signed Distance
$\mathrm{sd}(x_{t}):=\min_{\alpha\in\mathbb{R}}\quad\alpha\quad\mathrm{s.t.}\quad p\in\mathcal{O}_{1}^{x_{t}}(\alpha)\cap\mathcal{O}_{2}(\alpha).$ (7)
@@
Non-differentiability: For a $\{3,4\}$-dimensional polyhedron defined by $m_{\mathrm{sd}}$ inequality constraints, a vertex $w$ requires at least $\{3,4\}$ inequalities to become equalities. Every vertex can be attained by solving a linear system $Aw+b=0$, where $A\in\mathbb{R}^{(d+1)\times(d+1)}$, $b\in\mathbb{R}^{d+1}$. $A$ and $b$ depend on which inequalities we select to become equalities. If we traverse all possible selections, we can get all vertices.
@@
Given $x_{t}$, if there are only $d+1$ active constraints, there is only one optimal assignment, and we can get a unique assignment point $w$, so $\alpha$ is also unique, and the signed distance function $sd(x_{t})$ is smooth. However, when there are more than ${d+1}$ active constraints, the optimal assignment is not unique, and $sd(x_{t})$ is nonsmooth, because although $\alpha$ corresponding to these optimal assignments have the same value, they have different expressions and derivatives.
In Fig. 1, we show an example with two polygons and the corresponding 3-dimensional polyhedral LP space. There exist five feasible assignments, but only two values of $\alpha\in\{\alpha*,\alpha_{1}\}$. Clearly, $\alpha^{*}$ is the minimizer of Eq. 9, i.e. the actual signed distance. When $\alpha=\alpha^{*}$, the intersection point $p$ can be any point on the segment corresponding to an edge in the 3-dimensional space. Any point on this edge is a solution to the LP. If the ego rotates slightly, this edge will degenerate to either one of the vertices that correspond to $\alpha=\alpha^{*}$.
@@ III Collision Detection via Vertex Enumeration
To overcome this, we propose slots, placeholders which are dynamically filled with appropriate vertex values. Specifically, we let $N$ be the maximum number of vertices we wish to consider at any time. At each value of $x_{t}$, we enumerate all assignments, discard infeasible ones, and then sort the remaining $K$ assignments by $\mathrm{sd}_{k}(x_{t})$ value, and place $\mathrm{sd}_{k}(x_{t})$ expressions of the $N$ smallest ones into the slots as our signed distance constraints. If $K<N$, we copy the $\mathrm{sd}_{k}(x_{t})$ expression with the largest $sd_{k}(x_{t})$ value to fill the remaining slots, because the larger values are less important and correspond to inactive constraints. Choosing $N$ too small can adversely affect the algorithm success rate, but larger $N$ is more expensive to compute. The upper bound for meaningful $N$ is the most complicated situation when there is vertex to vertex contact. In the 2-dimensional case, $N\leq 4$ (two constraints from the ego and obstacle each), and more for higher dimensions, scaling with the sum of the maximum number of faces forming a vertex for the ego and obstacle.
The slot concept we introduce does not completely address the issue at hand. As $x_{t}$ is updated in the optimization, the particular vertices which fill each slot can change. Specifically, since each vertex is an assignment point, the assignment points which fill each slot can change. Although the assignment points are themselves smooth functions of $x_{t}$, this implies that the output value of each slot is not a smooth function of $x_{t}$. Nevertheless, we propose simply ignoring this fact when computing the derivative of each constraint slot (as needed by optimization solvers), and simply treating the derivative of the assignment point value (for the given value of $x_{t}$) as the derivative of the constraint slot.
The reformulation we have so far presented may seem counterintuitive. Initially, the signed-distance constraint appearing in Eq. 2 was a continuous but non-differentiable constraint. Using the slot formulation of Eq. 10, there are $K$ constraints which are also continuous but non-differentiable. However, there are two key benefits to this reformulation. The first benefit of this new formulation is representing derivative information at points of non-differentiability. In the formulation Eq. 10, non-differentiability occurs whenever there are multiple assignment points which could be placed into the constraint slots due to equivalent value. The only constraint slots which can ever be “active” are the first slot, and any other slots who are tied in value at a particular iteration. After some perturbation to $x_{t}$, the values will diverge and only the smaller of the two will remain relevant for the purposes of the optimization Eq. 10. It is plausible that the derivative information for that slot was incorrect before the perturbation, but if that was the case, then the correct derivative information must have been associated to one of the other tied slots. This enables the solver to essentially capture the subdifferential of the signed distance at these points of non-differentiability.
@@ IV-A Baselines
The first baseline method we compare to is the separating hyperplanes method [20, 21]. It is a common and efficient approach that utilizes the hyperplane separation theorem, which states there exists a separating hyperplane that separates two nonempty nonintersecting closed polyhedral objects. It is very fast for simple settings, where there is one ego and one obstacle. The second baseline method is [15], where the authors formulate the collision detection problem using a signed distance function as a convex optimization problem with conic constraints. However, this approach is flawed, because when inequality constraints are weakly active at the solution of the low-level problem, the implicit function theorem cannot be evoked, which our formulation avoids.
We compare our method to the baselines in terms of speed and reliability. Our metric for speed is the mean MCP solve time as solved by the PATH Solver. The reliability metric is more nuanced, we provide two metrics to represent reliability: (i) Success rate in terms of MCP solution achieved by PATH Solver represents the reliability of the method in successfully converging to a local optimum, (ii) and the mean cost of successes indicates how reliably the formulation achieves a low-cost local optimum. The success rate is the ratio of successful solutions to the total number of samples.
@@ IV-C Results
In this section we discuss large-scale 2-dimensional simulation results for our method and baselines on randomized obstacles and initial conditions for different problems. As we will show in this section, our method achieves higher success rates in settings where the ego needs to navigate through and pack into multiple obstacles, but it is slower than separating hyperplane formulation. We compare our method against two baseline methods for six different problems for randomly generated initial conditions and obstacles.
@@
We showcase the success rate of our method compared to baselines in Table I, where we use boldface to indicate the highest success rates. We see that the separating hyperplane approach is less reliable locally than our method for all but the simplest problem, simple packing. Our method achieves greater than $90\%$ success for all problems, beating the baselines in all settings except for the piano problem. We show the resulting cost for the successful solutions in Table II (failed MCP solutions are not included in this metric), where we use boldface to indicate the best cost. Table II tells us how successful each formulation is in terms of global success, such as being able to pass through a gap to reach the goal position. Our method outperforms the baselines in packing type problems, which means it packs more reliably compared to the baselines. In gap type problems, our method does not pass through the hole as often as [15] does. We believe this is because [15] tries to minimize contact with obstacles, while our method may generate a trajectory that maintains contact.
@@
Our nonsmooth signed distance enumeration method achieves superior accuracy solving complex collision detection problems, while remaining competitive in terms of performance. We have demonstrated that our formulation is more reliable compared to the separating hyperplane formulation, however it is slower. This is not surprising, as separating hyperplane formulation is quite efficient provided with a good initialization. Further, we show that our method is more reliable than a comparable but flawed approach in the literature for the problem of polygon collision detection, especially for packing problems where incorrect derivatives have a higher chance of becoming a problem for the solver.
@@ V CONCLUSION
In this paper we proposed a procedure for dealing with nonsmooth constraints by enumerating the values obtained at all vertices of the feasible region of signed distance between polyhedra. We then demonstrated our formulation on several challenging collision detection problems in 2 and 3-dimensional settings, and compared them to common existing formulations. We found that our method outperforms common baselines in terms of success rate. Our formulation is slower, but more reliable than separating hyperplanes.